OpenStudy (anonymous):

integrate -x(1+2x)^5 dx

OpenStudy (ash2326):

\[\int -x(1+2x)^5dx\] We can expand (1+2x)^5, but that will be tiring Do you know u substitution method ?

OpenStudy (anonymous):

how do i solve it by parts?

OpenStudy (ash2326):

Not by parts, do u know u-substitution? It'll make it easy

OpenStudy (ash2326):

@shaqadry Put 2x+1=t so \[2dx=dt\] \[dx=\frac{dt}{2}\] Do you get this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

wait but the question requires me to solve it by parts :-/

OpenStudy (ash2326):

Then, we will use this \[\int uv=u\int v-\int(du \int v)\]

OpenStudy (ash2326):

Can you try, apply it here?

OpenStudy (anonymous):

@ash2326 which one should i equate to 'u' ? which is the better choice?

OpenStudy (ash2326):

Equate x to u, it'll become 1 in the second part of by parts and then we find the integral easily @shaqadry

OpenStudy (anonymous):

okay thanks i will try to solve it!